OpenStudy (anonymous):

how do u simplify 8 over the radical of 2?? HELP ME!!

OpenStudy (anonymous):

\[8/\sqrt{2}=8\sqrt{2}/2=4\sqrt{2}\]

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

yw